# Towards More Realistic Extraction Attacks: An Adversarial Perspective

### v2

@@ 4 Experimental Setup
In this section, we outline our central experiment setup, to set the stage for our empirical study. Note that details about the setup for the case studies (§7) are delegated to their respective sections.
@@ 4.1 Models and Dataset
We use the Pythia suite (Biderman et al., 2023b) and OLMo models (Groeneveld et al., 2024) for all our experiments. We primarily focus on the Pythia suite, which contains decoder-only language models with the same architecture as EleutherAI’s GPT-Neo (Black et al., 2022), albeit different training, across various model sizes and with open source access to the complete training data and the intermediate checkpoints. Pythia models were trained using GPT-NeoX library (Andonian et al., 2023) on the Pile dataset (Gao et al., 2020) and have not undergone any form of instruction-tuning. The standard version of Pythia was trained over a single epoch of the Pile dataset, i.e., $\approx 143k$ steps with a batch size of $1024$, while the deduplicated version of Pythia was trained over $\approx 1.5$ epochs of the deduplicated Pile dataset, maintaining the same number of training steps as the standard version.
Pythia suite of models was developed with an emphasis on facilitating open-source investigation into the training dynamics of LLMs. As such, they offer access to (a) models of various sizes (we use model sizes: 1b, 1.4b, 2.8b, 6.9b, and 12b), (b) intermediate model checkpoints during training (a total of $154$ checkpoints, with $144$ of them equally spaced, i.e., at every $1k$ training steps), and (c) the complete training data order, which is the same for all model sizes. This level of accessibility and control over the training setup allows us to simulate the real-world availability of models across various sizes and with updating checkpoints over time.
To show the generalizability of our results, we also perform some additional experiments with OLMo models, another set of decoder-only language models. OLMo models were trained on the Dolma dataset Soldaini et al. (2024) and have also not undergone any form of instruction-tuning. These models also offer access to (a) intermediate model checkpoints during training, and (b) open-source access to the complete training data order.
5050505020020020020035035035035050050050050000%$1$%$2$%$3$%$4$%Composite Ext. Rate Prompt Length Extraction Rate (a) Extraction rate over Prompt Length
1b1.4b2.8b6.9b12b00%$0.5$%$1$%$1.5$%Composite Ext. Rate Model Size p (b) Extraction rate over Model Size p
100k110k120k130k140k00%$0.5$%$1$%$1.5$%$2$%Composite Ext. Rate Training Step (c) Extraction rate over Training Steps
10030050000%$1$%$2$%$3$% Prompt Length Extraction Rate (d) Extraction rate over Masking and Removal
BaselineMask ($5\%$)Mask ($10\%$)Remove ($5\%$)Remove ($10\%$)Composite Prefix00%$0.5$%$1$%Prefix g Extraction Rate (e) Extraction rate over Prefixes g No PrefixNumbersAlphanumericComposite
Figure 3: Extraction rates under prompt sensitivity and across multiple models for Pythia. (a) Increasing prompt length results in better extraction rates, with the composite extraction rate better than even at prompt length $500$. (b, c) We see similar trends for increasing model size and training steps, respectively. Specifically, we see the largest impact of the composite extraction rate across training steps, with the extraction rate increased 1.5×1.5$\times$1.5 × compared to any single checkpoint. (d) Randomly masking or removing tokens from the prompt severely hurts the extraction rate, highlighting the importance of prompt structure. (e) Adding a random prefix can also contribute to minor improvements in the composite extraction rate.
@@ 4.2 Evaluation Methodology
We now describe our approach to the design and evaluation of extraction attacks. Similar to Carlini et al. (2023), we sample a representative portion of the dataset for analyzing the performance of our extraction attacks. More specifically, we uniformly sample $100,000$ sequences from the first $100k$ steps (batches) of the training data for Pythia. This sampling strategy is important because we choose model checkpoints for evaluation starting at step $100k$, which ensures that every sentence evaluated for memorization has been seen by each checkpoint under consideration, as illustrated in Figure 2. We use the same approach for OLMo, with the training step $300k$ being the cut-off point.
Each sequence sampled is exactly $2049$ tokens. For our analysis, we employ a consistent method of partitioning each sequence into a prompt and completion at the midpoint, i.e., $1024$ tokens. Formally, for a sentence $s_{1:2049}$, prompt length $l_{p}$, and completion length $l_{x}$, the example $[{\mathbf{p}}\parallel{\mathbf{x}}]$ is defined as ${\mathbf{p}}=s_{1024-l_{p}:1024}$ and ${\mathbf{x}}=s_{1024:1024+l_{x}}$. This partitioning allows us to systematically vary the prompt length and design while comparing the same completion, and vice-versa.
For the Pythia suite, unless otherwise specified, we use a prompt length of $l_{p}=50$, a completion length of $l_{x}=50$, the Pythia-6.9b model, and the 140k training step checkpoint, evaluating the extraction attacks using verbatim match. We use the same default setup for OLMo, with the OLMo-7b model, and the 500k training step checkpoint.
@@ 5.1 Prompt Sensitivity
Prompt Length. Prompt length is a commonly studied parameter in extraction attacks, and it has been shown that longer prompts lead to better extraction (Carlini et al., 2023). This is intuitive, as conditioning the model with more text from training would increase the likelihood of extraction. We will now show that the composite extraction rate (Definition 3.1) across varying prompt lengths exceeds the extraction rate at even the largest prompt length. As illustrated in Figure 3(a), the extraction rate increases with longer prompts. However, the composite extraction rate is noticeably higher than any single prompt length, including the longest at 500 tokens. This suggests that certain information extractable with shorter prompts remains elusive even with the longest prompt. Consequently, an adversary can exploit this churn across the prompt length to extract more information. We see similar trends for OLMo in Figure 4.
Prompt Length Training Steps 00%$0.5$%$1$%$1.5$%$2$%$2.5$% Extraction Rate
Prompt Length $50$Training Step $300k$Prompt Length $500$Training Step $500k$Composite All
@@
Figure 4: Composite extraction attack results across 10 prompt lengths (same as Pythia) and 11 training steps (equidistant between $300k$ and $500k$), compared against isolated setups, for OLMo.
@@
Prompt Structure. Next, we explore the structure of prompts to identify what makes a prompt potent and where churn can emerge. We introduce noise into the prompts by masking and removing random tokens; results are collected in Figure 3(d). Despite introducing only a small amount of noise, we observe a significant drop in extraction rates. This indicates that the contiguous prompt from the training data is crucial for extracting information, and any disruption inside this prompt can significantly hurt its capabilities. Yet, we do see minute churn in extraction trends, which further highlights how an adversary can exploit repeated prompting to extract more information, even with seemingly unintuitive changes like masking or removing random tokens.
We next add noise as a prefix in the form of random numeric and alphanumeric strings; results are collected in Figure 3(e). Interestingly, the performance degradation with a noisy prefix is less severe than with noise within the prompt. More importantly, we observed a higher composite extraction rate. This suggests that adding a noisy prefix can also help extract unique information that was previously inaccessible, and further highlights how an adversary can exploit repeated prompting.
@@
Note that the churn in our prompt design trends highlights the increased extraction risks without access to new information. For instance, if an adversary has access to the prompt of length 500 tokens, they can expand the attack surface and thereby the extraction rate simply by removing parts of the prompt, adding noise, etc.—without needing any additional knowledge. One might argue that as the number of prompt variations increases, every sentence could become extractable. However, that is not true; not all sentences are extractable. Yin et al. (2024) showed that knowledge not present in an LLM will not be extractable even after prompt optimization, while Schwarzschild et al. (2024) also showed similar trends when attempting to extract a given completion. Consequently, prompting an LLM to regurgitate certain sentences, even with various prompt modifications, demonstrates a genuine extraction risk and underscores the extent of memorization in LLMs (Carlini et al., 2021).
@@ 5.2 Multiple Checkpoints
Model Size. The model size has long been known to influence learning trends, and our results in Figure 3(b) reflect this phenomenon. We find that larger models tend to memorize more information, which makes them more vulnerable to extraction attacks. However, our results also indicate that the composite extraction rate is higher than the extraction rate of any single model, highlighting the churn present in these trends. Biderman et al. (2023a) also conducted an empirical study on the overlap between memorized data across model sizes and found that up to $10\%$ of the data memorized by smaller models is not memorized by larger models. Combining our insights with existing literature, it’s clear that releasing models in different sizes increases the extraction risks.
@@
PilePile-Deduped00%$1$%$2$%$3$% Dataset Extraction Rate (b) gComposite Attacks and Deduplicationg
[…]
Figure 5: Towards more realistic extraction rates by combining various churn trends and with approximate matching. (a) Combining two axes at a time, we see a monotonically increasing trend in extraction rates as we gain more points of access to the underlying dataset, highlighting the growing power of the adversary. (b) Combining multiple axes of attack results in a significant increase in extraction rate for both standard and deduplicated setups, with the composite extraction rate for the deduplicated setup the same as the highest single setting rate for the standard setup. (c) Various similarity metrics have distinct trends as we decrease the threshold value and allow for looser approximations, thus the choice is context-driven.
Model Updates. We also analyze model updates over time using intermediate checkpoints in Figure 3(c), where we observe the most significant churn in our study. Unsurprisingly, attacking models at later stages of training is more successful, as seen in the literature (Tirumala et al., 2022; Biderman et al., 2023a; Jagielski et al., 2023). But remarkably, the churn here is significantly powerful and by exploiting composability across intermediate checkpoints, an adversary can increase their extraction rate by more than 1.5×1.5$\times$1.5 ×. We also see similar results for OLMo in Figure 4. This underscores the impact of stochasticity in model training on extraction attacks and reveals that regular model updates, typically considered beneficial in the current LLM ecosystem, create a powerful adversary.
6 Towards Realistic Extraction Attacks
With a better understanding of the trends across various setups, we now evaluate a more realistic measure of leakage in extraction attacks, by investigating (a) composability in churn, (b) challenges in evaluation, and (c) effects of deduplication.
@@ 6.1 Combining Multiple Axes of Churn
In the previous section, we saw how churn can impact individual axes of variability, such as prompt sensitivity, model size, and intermediate checkpoints. However, a real-world adversary can take advantage of all these factors simultaneously, thus significantly increasing their extraction rates. We start by analyzing two axes at a time, as shown in Figure 5(a). For all pairs of variability, the overall composite extraction rate (bottom right) is 23×2-3$\times$2 - 3 × higher than the base setup (top left) and 1.52×1.5-2$\times$1.5 - 2 × higher than the composite extraction rates along one axis (top right and bottom left). Furthermore, when all three axes are combined, depicted in Figure 5(b), the extraction rates grow even higher, albeit with diminishing gains. Thus, we show that a real-world adversary can extract far more training data than has been previously seen in the literature.
5005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 000000.20.20.20.20.40.40.40.4 Train Dataset Size $p$-value ($\downarrow$) (a) Dataset Inference
10010010010015015015015020020020020025025025025030030030030000%$5$%$10$%$15$% ROUGE-L Length Extraction Rate ($\uparrow$) (b) Copyright Infringement
Best Single SetupComposite Prompt LengthsComposite Model SizesComposite Training Steps
Figure 6: (a) $p$-value for dataset inference (lower is better) across different dataset sizes. The results show the importance of exploiting prompt sensitivity and the significant improvement under the composite setup across different prompt lengths. (b) Extraction rate for different ROUGE-L length thresholds, marking potential copyright violations generated by the LLM. Extraction rates with composite setups are consistently higher than the single setup, highlighting the impact of multi-faceted access to the data.
@@ 6.2 Approximate Matching
As discussed in §3.2, evaluating extraction attacks under verbatim match can underestimate the true risk of extraction. To address this gap, we introduced approximate composite discoverable memorization (Definition 3.2), and will now analyze various similarity metrics $S$ to examine their behaviour under changing $\delta$, reported in Figure 5(c). Solely for this discussion, we increase the completion length $l_{x}=500$, to allow for meaningful extraction even with approximate matching.
Our results reveal intriguing trends. First, we analyze evaluations based on the Levenshtein ratio metric and observe that even the threshold of $\delta=0.95$ doubles the extraction attack rate compared to a verbatim match. This threshold signifies a minimum $95\%$ overlap between generated and original text. Even under such a strict threshold, the doubling of the extraction rate underscores the significant underestimation of extraction risks when relying solely on verbatim matches. As $\delta$ decreases, however, the extraction rate increases exponentially, as the Levenshtein ratio becomes less reliable under looser constraints. We also see similar trends for ROUGE-L scores.
Transitioning to other similarity metrics — longest common substring (LCS), Hamming distance, and $n$-gram matching — we find that even lower values of similarity ($\delta$) can contribute meaningfully to extraction attacks. We observe patterns of rising extraction rates similar to what we saw earlier with the Levenshtein distance. The diverse trends underscore the choice of approximation metric as highly context-dependent. A more thorough examination of which metrics best serve particular applications is left for future work.

### v3

@@ 4 Experimental Setup
In this section, we outline our central experiment setup to set the stage for our empirical study. Further details about the setup for the case studies (§7) are delegated to their respective sections.
@@ 4.1 Models and Dataset
We use the Pythia suite (Biderman et al., 2023b) and OLMo models (Groeneveld et al., 2024) for all our experiments. We focus on the Pythia models throughout the paper, while also providing complementary results on OLMo models to show the generalizability of our analysis. Pythia suite offer access to (a) models of various sizes (1b, 1.4b, 2.8b, 6.9b, and 12b), (b) intermediate checkpoints during training (a total of $154$ checkpoints, with $144$ of them equally spaced, i.e., at every $1k$ training steps), and (c) the training data (Pile dataset (Gao et al., 2020)) as well as the training order, which is the same for all model sizes. This level of accessibility and control over the training setup allows us to simulate the real-world availability of models across sizes and different checkpoints over time.
OLMo models were trained on the Dolma dataset Soldaini et al. (2024) and also offer access to (a) intermediate model checkpoints during training, and (b) the complete training data order.
@@ 4.2 Evaluation Methodology
We now describe our evaluation methodology. Similar to Carlini et al. (2023), we sample a representative portion of the training data for analyzing the performance of our extraction attacks. More specifically, we uniformly sample $100,000$ sequences (prompts) from the first $100k$ steps (batches) of the training data for Pythia. This sampling strategy is important because we choose model checkpoints for evaluation starting at step $100k$, which ensures that every sentence evaluated for memorization has been seen by each checkpoint under consideration (as illustrated in Figure 3).
Each sequence sampled is exactly $2049$ tokens. For our analysis, we employ a consistent method of partitioning each sequence into a prompt and completion at the midpoint, i.e., $1024$ tokens. Formally, for a sentence $s_{1:2049}$, prompt length $l_{p}$, and completion length $l_{x}$, the example ${\mathbf{p}}\parallel{\mathbf{x}}]$ is defined as ${\mathbf{p}}=s_{1024-l_{p}:1024}$ and ${\mathbf{x}}=s_{1024:1024+l_{x}}$. This partitioning allows us to systematically vary the prompt length and design while comparing the same completion, and vice versa. We use the same approach for OLMo, with the training step $300k$ (instead of $100k$) being the cut-off point.
For the Pythia suite, unless otherwise specified, we use a prompt length of $l_{p}=50$, a completion length of $l_{x}=50$, the Pythia-6.9b model, and the 140k training step checkpoint, evaluating the extraction attacks using verbatim match. For OLMo, we use the OLMo-7b model and the 500k training step checkpoint as defaults, while the rest of the configuration is the same as Pythia.
@@ 5 Churn in Extraction Trends
50502002003503505005000%11%22%33%44%Composite Ext. Rate Prompt Length Extraction Rate (a) Across Prompt Length
1b1.4b2.8b6.9b12b0%0.50.5%11%1.51.5%22%Composite Ext. Rate Model Size p (b) Across Model Size p
100k110k120k130k140k0%0.50.5%11%1.51.5%22%Composite Ext. Rate Training Step (c) Across Training Steps
1003005000%11%22%33% Prompt Length Extraction Rate (d) Masking and Removal
BaselineMask ($5\%$)Mask ($10\%$)Remove ($5\%$)Remove ($10\%$)Composite Prefix0%0.50.5%11%Prefix g Extraction Rate (e) Random Prefixes g No PrefixNumbersAlphanumericComposite
Figure 4: Extraction rates under prompt sensitivity and across models for Pythia. (a) Increasing prompt length improves extraction, with the composite extraction rate better than even at prompt length $500$. (b, c) Increasing model size and training steps show similar trends, with the largest impact of the composite extraction rate seen in training steps, increasing the extraction rate $1.5\times$ compared to a single checkpoint. (d) Randomly masking or removing tokens severely hurts the extraction rate, highlighting the importance of prompt structure. (e) Adding a random prefix can also contribute to the composite extraction rate.
@@ 5.1 Prompt Sensitivity
Prompt Length. Prompt length is a commonly studied parameter in extraction attacks, and it has been shown that longer prompts lead to better extraction (Carlini et al., 2023). This is intuitive, as conditioning the model with more text from training would increase the likelihood of extraction.
However, we show in Figure 4(a) that the composite extraction rate (Definition 3.1) across varying prompt lengths is noticeably higher than the extraction rate at even the largest prompt length at 500 tokens. Thus, certain information extractable with shorter prompts remains elusive even with the longest prompt, due to prompt sensitivity in LLMs, as discussed in §3.1. Consequently, an adversary can exploit this churn across the prompt length to extract more information. We see similar trends for OLMo in Figure 5.
Prompt Length Training Steps 0%0.50.5%11%1.51.5%22%2.52.5% Extraction Rate
Prompt Length 5050Prompt Length 500500Training Step 300k300kTraining Step 500k500kComposite All
@@
Figure 5: Composite extraction attack results across 10 prompt lengths (same as Pythia) and $1$ training steps (equidistant between $300k$ and $500k$), compared against isolated setups, for OLMo.
Figure 6: Real examples of churn. (Left) Prompts of different lengths can contribute uniquely to the model extraction, and longer contexts aren’t always better. (Right) Similarly, different model sizes contribute uniquely to the composite extraction rate, showing the importance of churn across model size.
@@
Prompt Structure. Next, we explore the structure of prompts to identify where churn can emerge. We introduce noise into the prompts by masking and removing random tokens (Figure 4(d)); and as a prefix in the form of random numeric and alphanumeric strings (Figure 4(e)). Despite introducing only a small amount of noise, we observe a significant drop in extraction rates. This indicates that the contiguous prompt from the training data is crucial for extracting information, and any disruption to this prompt can significantly hurt its capabilities.
Yet, we do see churn in extraction trends, with a larger impact of the noisy prefixes. This further highlights how an adversary can exploit repeated prompting, even with seemingly unintuitive changes like masking or removing random tokens and adding a random prefix to the input prompt.
Note that the churn trends under prompt sensitivity, both for prompt structure and prompt length, highlight the increased extraction risks without access to new information. For instance, if an adversary has access to the prompt of length 500 tokens, they can expand the attack surface and thereby the extraction rate simply by removing parts of the prompt (prompt length composite attack from Figure 4(a)), adding noise (prompt structure composite attack from Figure 4(d, e)), etc., without needing any additional knowledge.
@@
One might argue that as the number of prompt variations increases, every sentence could become extractable. However, that is not true; not all sentences are extractable. Yin et al. (2024) showed that knowledge not present in an LLM will not be extractable even after prompt optimization, while Schwarzschild et al. (2024) also showed similar trends when attempting to extract a given completion. Consequently, prompting an LLM to regurgitate certain sentences, even with various prompt modifications, demonstrates a genuine extraction risk and underscores the extent of memorization in LLMs (Carlini et al., 2021).
123450%0.20.2%0.40.4%0.60.6% Number of Models Extraction Rate (a) Recurrence of Extractions across Model Size
1b1.4b2.8b6.9b12b0%0.10.1%0.20.2% Model Size (b) Unique Extractions across Model Size
1b1.4b2.8b6.9b12b0%0.50.5%11%1.51.5% Model Size (c) Drop-One Extractions across Model Size
Figure 7: Granular trends of churn across model size for Pythia. (a) Number of models that successfully extract the same information, i.e., whether only 1 out of 5 models extracts a given sample, or whether 2, 3, 4, or all 5 do. (The x-axis here indicates the exact number of models that extract the same samples.) Even though the majority of samples are extractable by all 5 models, a significant amount of extractions are unique to a subset, or even a single model. (b) While larger models contribute more unique extractions, each model, regardless of size, provides some unique extractions. (c) Composite extraction rates after dropping a single model (x-axis: dropped model, y-axis: resulting extraction rate) show that extraction rates remain stable even if we remove the contributions of the biggest model.
@@ 5.2 Multiple Checkpoints
Model Size. The model size has long been known to influence learning trends, and our results in Figure 4(b) reflect this phenomenon. Larger models tend to memorize more information, which makes them more vulnerable to extraction attacks. However, our results also indicate that the composite extraction rate is higher than the extraction rate of any single model, highlighting the churn in these trends. Biderman et al. (2023a) also conducted an empirical study on the overlap between memorized data across model sizes and found that up to $10\%$ data memorized by smaller models is not memorized by larger models. Combining our insights with existing literature, it’s clear that releasing models in different sizes increases the extraction risks.
Model Updates. We also analyze model updates over time using intermediate checkpoints in Figure 4(c), where we observe the most significant churn in our study. Unsurprisingly, attacking models at later stages of training is more successful, as seen in the literature (Tirumala et al., 2022; Biderman et al., 2023a; Jagielski et al., 2023). But remarkably, the churn here is significantly powerful and by exploiting composability, an adversary can increase their extraction rate by more than $1.5\times$. This underscores the impact of stochasticity in model training on extraction attacks and reveals that regular model updates, typically considered beneficial in the LLM ecosystem, create a powerful adversary. We also see similar results for OLMo in Figure 5.
We provide some examples of extractions from the Pile dataset that show regressive trends, i.e., successful extraction using weaker setups, and highlight the value of churn, in Figure 6.
5.3 Unique Extractions
Next, we study the contribution of each setup to the composite extraction rate. We focus on trends across model size, while providing results for other axes in the appendix (§A). We find that while a majority of extracted samples are extractable with all 5 model sizes, there is a significant portion of extractions unique to a few (or even just one) model(s) (Figure 7(a)). Examining the distribution of extractions unique to individual models (Figure 7(b)), we observe that Pythia-6.9b produces a substantial number of unique extractions not found even in the larger Pythia-12b. Smaller models, such as Pythia-1b, 1.4b, and 2.8b, also contribute non-trivially.
Finally, we study composite extraction rates when all models except one are used (Figure 7(c)), to quantify the contribution of each model. We find that even after dropping the largest model, the remaining models achieve high composite extraction rates, indicating that no single model is essential for strong extraction. Given that attacking the most vulnerable model can be expensive, this shows that an adversary can take advantage of the churn to instead perform a composite attack on less vulnerable but cheaper models (more details in §6.1).
6 Towards Realistic Extraction Attacks
With a better understanding of the churn, we now evaluate a more realistic measure of leakage in extraction attacks, by investigating (a) composability across multiple axes, (b) cost of composite attacks, (c) approximate matching, and (d) deduplication.
@@ 6 Towards Realistic Extraction Attacks
PilePile-Deduped0%11%22%33% Dataset Extraction Rate (b) gComposite Attacks and Deduplicationg
Prompt Length 100100Training Step 100k100kModel Size 1.4b1.4bPrompt Length 300300Training Step 140k140kModel Size 12b12bComposite All 0.10.10.40.40.70.7110%22%44% Similarity Threshold ($\delta$) Extraction Rate (at $l_{x}=500$) (c) Comparing Similarity Metrics VerbatimLevenshteinLCSnn-gramHammingROUGE-L
Figure 8: Towards more realistic extraction by combining various churn trends and with approximate matching. (a) Combining two axes at a time, we see a monotonically increasing trend in extraction rates as we gain more points of access to the underlying data, highlighting the growing power of the adversary. (b) Combining all axes of attack results in a significant increase in extraction rate for both standard and deduplicated setups. (c) Various similarity metrics have distinct trends as we decrease the threshold and allow for looser approximations, thus the choice of similarity metric depends on the context of the attack.
@@ 6.1 Combining Multiple Axes of Churn
In the previous section, we saw how churn can impact individual axes of variability. However, a real-world adversary can take advantage of all factors simultaneously, thus significantly increasing their extraction rates. We start by analyzing two axes at a time in Figure 8(a). For all pairs of variability, the overall composite extraction rate (bottom right) is $2-3\times$ higher than the base setup (top left) and $1.5-2\times$ higher than the composite extraction rates along one axis (top right and bottom left). Furthermore, when all three axes are combined, depicted in Figure 8(b), the extraction rates grow even higher, albeit with diminishing gains. Thus, a real-world adversary can extract far more training data than shown in existing literature.
6.2 Computation Cost of Composite Attacks
We now study the computational cost of performing composite extraction attacks. We define our cost in units relative to attacking the Pythia-1.4b model with a prompt length of 100 (the cheapest setup to attack in our multiple axes setting). Cost scales linearly with model size (i.e., attacking the Pythia-2.8b model with prompt length 100 costs 2 units) and quadratically with prompt length (i.e., attacking the Pythia-1.4b model with prompt length 200 costs 4 units), while remaining constant across different checkpoints. The resulting extraction trends for all axes of churn combined under varying cost budgets are shown in Figure 9.
0101020203030404050501.51.5%22%2.52.5%33% Computation Cost Units Extraction Rate
Single GreedyCombined GreedyCombined Best
Figure 9: Various strategies of utilizing the adversarial budget and the resulting extraction rates.
We study three different strategies for utilizing the given resources. First, we consider an adversary that greedily attacks the most expensive setup, i.e, prioritizing the largest prompt length, model size, and latest checkpoint, while leaving any leftover resources unused. The extraction rate stays flat until the budget can support a more expensive setup, causing sharp jumps and creating a staircase pattern. This strategy mirrors prior work, where each setup is evaluated in isolation. Next, we extend the previous strategy to utilize the unused resources, and instead of attacking only a single setup, the adversary now selects multiple setups, still greedily choosing to fit the remaining budget.
Finally, we define the third strategy, which is the most effective: searching all combinations to maximize the composite extraction rate, rather than greedily targeting the most vulnerable setups. This consistently outperforms the staircase trend even at the jumps, i.e., even when the greedy strategy utilizes all available budget, showing that combining extractions from less vulnerable setups can exceed the returns of attacking the most expensive setup.
@@ 6.3 Approximate Matching
As discussed in §3.2, evaluating extraction attacks under verbatim match can underestimate the true risk of extraction. Here, we analyze approximate composite discoverable memorization (Definition 3.2) across various similarity metrics $S$ to examine their behaviour under changing $\delta$ in Figure 8(c). Solely for this discussion, we increase the completion length $l_{x}=500$, to allow for meaningful extraction even with approximate matching.
Our results reveal intriguing trends. First, we study evaluations based on the Levenshtein ratio metric and observe that even the strict threshold of $\delta=0.95$ doubles the extraction attack rate compared to a verbatim match. This threshold signifies a minimum $95\%$ overlap between generated and original text. Clearly, we miss out on a considerable amount of leaked information by relying only on verbatim matches. As $\delta$ decreases, the extraction rate increases exponentially, as the Levenshtein ratio becomes less reliable under looser constraints. We also see similar trends for ROUGE-L scores.
For similarity metrics like longest common substring (LCS), Hamming distance, and $n$-gram matching, even lower values of similarity ($\delta$) can contribute meaningfully to extraction attacks. Unsurprisingly, we observe patterns of increasing extraction rates as before, albeit slower. The diverse trends underscore the choice of the approximation metric as highly context-dependent. A more thorough examination of which metrics best serve particular applications is left for future work.